### v2

# PEAN: A Diffusion-Based Prior-Enhanced Attention Network for Scene Text Image Super-Resolution

Abstract.
Scene text image super-resolution (STISR) aims at simultaneously increasing the resolution and readability of low-resolution scene text images, thus boosting the performance of the downstream recognition task. Two factors in scene text images, visual structure and semantic information, affect the recognition performance significantly. To mitigate the effects from these factors, this paper proposes a Prior-Enhanced Attention Network (PEAN). Specifically, an attention-based modulation module is leveraged to understand scene text images by neatly perceiving the local and global dependence of images, despite the shape of the text. Meanwhile, a diffusion-based module is developed to enhance the text prior, hence offering better guidance for the SR network to generate SR images with higher semantic accuracy. Additionally, a multi-task learning paradigm is employed to optimize the network, enabling the model to generate legible SR images. As a result, PEAN establishes new SOTA results on the TextZoom benchmark. Experiments are also conducted to analyze the importance of the enhanced text prior as a means of improving the performance of the SR network. Code will be made available at https://github.com/jdfxzzy/PEAN.
Scene Text Image, Super-Resolution, Vision Backbone, Diffusion Models
ccs: Computing methodologies Reconstruction
1. Introduction
Scene text recognition (STR) focuses on extracting text from images, which has been widely applied in automatic driving (Reddy et al., 2020), intelligent transportation (Al-Shemarry et al., 2023), etc. However, in real-world applications, a variety of reasons result in captured images being low-resolution (LR), such as the quality of the lens, motion blur, and shaking when capturing photos, leading to blurred text in images. To better read text from such images, researchers formulate the STISR task to reconstruct missing text details in LR images, as a pre-processing step for STR.
Figure 1. Comparison between previous text prior-based STISR methods (row (b, c)) and PEAN. The incorporation of AMM enables PEAN to restore the visual structure of lengthy text in images. However, its performance is limited by the absence of semantic information (row (d)). The introduction of TP-LR partially addresses this limitation, yet its efficacy remains inadequate, leading to several failure cases (row (e)). Considering that TP-HR is a robust alternative, we conduct an exploratory experiment by substituting TP-HR with TP-LR, resulting in superior performance (row (f)). This inspires us to design a module for enhancing the TP-LR so as to obtain the ETP, which demonstrates comparable effectiveness to TP-HR in guiding the SR process (row (g)).
For scene text images, two crucial factors determine whether they could be correctly recognized, i.e., visual structure and semantic information (Fang et al., 2021a; Yu et al., 2020). Early attempts at STISR concentrate on adequately recovering the visual structure of LR scene text images (Fang et al., 2021b; Wang et al., 2020; Zhao et al., 2021). Composed of several CNN-BiLSTM layers, these methods can learn from paired LR-HR images to improve the resolution and readability of scene text images simultaneously. However, the performance is limited due to the fact that they ignore the semantic information of scene text images. This factor has been utilized in recent advancements. These works observe that semantic information plays an important role in guiding the restoration of correct visual structure, and propose numerous text prior-based methods (Guo et al., 2023; Ma et al., 2023, 2022; Zhao et al., 2022). That is, the text prior, generated by pre-trained STR models, is leveraged to facilitate the SR process (Guo et al., 2023; Ma et al., 2023, 2022; Zhao et al., 2022), thereby generating correct characters of text in SR images.
Despite improved performance achieved by these approaches, the dominance of visual structure and semantic information persists, as two critical issues in previous studies remain unresolved. Firstly, previous STISR methods (Chen et al., 2021, 2022; Ma et al., 2022; Wang et al., 2020; Zhao et al., 2022) rely on Sequential Residual Blocks (SRB) to extract visual features. This module, containing several CNN-BiLSTM layers, has difficulty in restoring the complete visual structure of images containing long or deformed text string due to its inherent demerits, i.e., the performance bottleneck of capturing long-range dependence (Dosovitskiy et al., 2020; Raghu et al., 2021). Secondly, the introduction of the primary text prior, originating from the interference of low-quality images on recognizers, prevents the SR network from generating images that contain correct semantic information. Recently, C3-STISR (Zhao et al., 2022) has employed a language model (Fang et al., 2021a) into STISR, utilizing its learned linguistic knowledge to rectify the text prior. Although the rectified prior demonstrates some effectiveness, it lacks sufficient strength in guiding the SR network to produce images with high semantic accuracy.
We propose a Prior-Enhanced Attention Network (PEAN) to tackle issues caused by the two factors. To begin with, an Attention-based Modulation Module (AMM) is proposed to substitute the SRB, endowing the network with a larger receptive field to images, thereby restoring the visual structure of images with text in various shapes and lengths (shown in Figure 1(d)). Horizontal and vertical strip-wise attention mechanisms (Dong et al., 2022; Huang et al., 2023; Tsai et al., 2022) are employed in AMM. Among them, the horizontal attention mechanism can capture the dependence between characters while the vertical attention mechanism can capture the structural information within a character (Zhao et al., 2021). However, the lack of semantic information limits the capability of such model. As demonstrated by previous works (Dong et al., 2016; Dahl et al., 2017), leveraging strong prior information to restrict the solution space plays a vital role in SR problems. Notably, the text prior derived from high-resolution (HR) images is a robust choice for STISR, in view of the high recognition accuracy of HR images. Consequently, we conduct an exploratory experiment wherein we substitute the text prior from LR images (TP-LR) with the text prior from HR images (TP-HR) within such model, yielding superior outcomes (see Figures 1(e) and (f) for comparison, details can be found in § 3). This inspires the design of a module for enhancing the primary text prior, resulting in the creation of the Enhanced Text Prior (ETP), which is comparable in effectiveness to TP-HR (shown in Figures 1(f) and (g)). The ETP provides valuable guidance to the SR network, promoting the generation of SR images with high semantic accuracy. Given the remarkable performance of diffusion models (Ho et al., 2020; Song et al., 2021), we propose a diffusion-based Text Prior Enhancement Module (TPEM) to obtain the ETP owing to their ability to map complex distributions (Xia et al., 2023). In addition, considering that the goal of STISR is to increase the resolution and readability of LR scene text images, we adopt the Multi-Task Learning (MTL) paradigm in the training phase, where the image restoration task aims at generating high-quality SR images, and the text recognition task stimulates the model to generate more readable SR results. In a nutshell, main contributions of our work are three-fold:
  • We devise an AMM containing horizontal and vertical attention mechanisms to model the long-range dependence in scene text images, thereby recovering the visual structure of images with long or deformed text.
  • A diffusion-based TPEM is further proposed to enhance the primary text prior. The resulting ETP guides the SR network to generate SR images with improved semantic accuracy.
  • Empirical studies show that PEAN attains the SOTA performance on the TextZoom (Wang et al., 2020) benchmark. We also conduct experiments to explore the reasons behind the performance improvement of the SR network.
2. Related Work
In this section, we give a brief review of recent progress in scene text image super-resolution, scene text recognition and diffusion models, which are relevant to our work.
2.1. Scene Text Image Super-Resolution
Scene text image super-resolution (STISR) has received surging attention in the computer vision community. Different from the classic single image super-resolution (SISR) task, STISR aims at increasing the resolution and legibility of scene text images simultaneously (Zhu et al., 2023b), serving as a pre-processing method for the downstream recognition task.
The milestone works in STISR are the TextZoom benchmark and the TSRN model (Wang et al., 2020), which promote the development of follow-up approaches. We roughly classify them into two categories. One category of methods focuses on recovering the visual structure of LR scene text images. Among them, TSRN (Wang et al., 2020) and PCAN (Zhao et al., 2021) use several CNN-BiLSTM blocks to complete the SR process. TSAN (Zhu et al., 2023a) adopts a gradient-based graph attention method to extract more effective representations for STISR. Another category considers the semantic information as the text prior to guide the SR process. In this category, TPGSR (Ma et al., 2023) and TATT (Ma et al., 2022) utilize the pre-trained recognizer to generate the text prior from LR images, boosting the model to generate SR images with correct text. C3-STISR (Zhao et al., 2022) employs a language model (Fang et al., 2021a) to rectify the text prior and uses triple clues to realize STISR.
Figure 2. Overview of the architecture of our proposed Prior-Enhanced Attention Network (PEAN).
2.2. Scene Text Recognition
Scene text recognition (STR) aims at reading text contents from natural images. The pioneering work CRNN (Shi et al., 2016a) uses the CNN-BiLSTM framework and the CTC (Graves et al., 2006) loss to perform STR for the first time. ASTER (Shi et al., 2018) further exploits the Thin-Plate Spline (TPS) transformation (Jaderberg et al., 2015) to understand scene text images with deformed or irregular layouts. The concurrent work MORAN (Luo et al., 2019) also handles these cases via a multi-object rectification network.
Recently, language models have been integrated into STR models and the fusion of vision and language features shows a great potential to improve the scene text understanding. SRN (Yu et al., 2020) employs an autoregressive language model to rectify the recognition results generated by visual features. Additionally, ABINet (Fang et al., 2021a) shows that masked language models (Devlin et al., 2019), capable of providing bidirectional representations, constitute another effective option for rectification. Furthermore, PARSeq (Bautista and Atienza, 2022) creatively adopts an internal language model as a spell-checker, eliminating the need for the pre-training process in ABINet (Fang et al., 2021a).
2.3. Diffusion Models
In computer vision, diffusion models (Ho et al., 2020) emerge as robust probabilistic generative models, facilitating tasks like image synthesis (Gao et al., 2023b; Rombach et al., 2022), text-to-image synthesis (Nichol et al., 2022; Saharia et al., 2022), image restoration (Fei et al., 2023; Xia et al., 2023) and image inpainting (Zhu et al., 2024a) through the iterative diffusion of information among pixels. Recently, diffusion models have also been employed in the super-resolution task. SR3 (Saharia et al., 2023) is the pioneering work that applies diffusion models to SISR. TextDiff (Liu et al., 2023) represents the initial attempt at a diffusion-based model designed specifically for STISR, focusing on enhancing the visual structure of text within images by refining their contours for a more natural appearance.
In contrast to existing methods, the proposed PEAN uses the diffusion-based TPEM to provide the SR network with enhanced semantic guidance, further resulting in SR images with heightened semantic accuracy.
3. Methodology
This section first gives an overview of the Prior-Enhanced Attention Network (PEAN). Then we present the proposed Text Prior Enhancement Module (TPEM), Attention-based Modulation Module (AMM) and the Multi-Task Learning (MTL) paradigm.
3.1. Overall Architecture
The pipeline of our proposed PEAN is shown in Figure 2. Given one LR image $I^{\mathrm{lr}}\in\mathbb{R}^{H\times W\times C}$, the Text Prior Generator (TPG) outputs the recognition probability sequence as the primary text prior $P^{\mathrm{l}}$. Then, the diffusion-based TPEM refines it to obtain the ETP denoted as $P^{\mathrm{e}}$, which can assist the SR network to generate SR images with improved semantic accuracy. Concurrently, a convolutional layer is adopted to extract the shallow visual feature $F^{\mathrm{s}}$ from $I^{\mathrm{lr}}$, which is then aligned with the refined text prior by a Feature Alignment Module (FAM). Then an AMM with $N$ blocks is introduced to mine the internal dependence between characters in the image, thereby facilitating the SR process. For the $i^{\text{th}}$ block of AMM (i.e., $B_{i}$), its output $F_{i}^{\mathrm{o}}$ is firstly concatenated with the aligned feature (i.e., $F^{\mathrm{a}}$) in the channel dimension to get $F_{i+1}^{\mathrm{c}}$. The fusion feature is then sent into $B_{i+1}$ for further processing. Finally, a Super-Resolution Module (SRM) containing several convolutional and batch normalization layers, receives $F_{N}^{\mathrm{o}}$ as input and utilizes a PixelShuffle (Shi et al., 2016b) operation to generate the SR image $I^{\mathrm{sr}}\in\mathbb{R}^{2H\times 2W\times C}$. Notably, in the training phase, $F_{N}^{\mathrm{o}}$ is also sent into an Auxiliary Recognition Module (ARM), which outputs the recognition probability sequence of the SR image. The outputs of TPEM, SRM and ARM enable the optimization of the model in an Multi-Task Learning (MTL) paradigm, steering the model to generate plausible and readable SR images.
Figure 3. Overview of the architecture of the FAM and the strip-wise attention mechanism inside LAM and GAM.
3.2. Text Prior Enhancement Module
As demonstrated by previous works (Dahl et al., 2017; Dong et al., 2016), strong prior information plays a pivotal role in solving SR problems, while the primary text prior applied in previous works is not powerful enough because it originates from LR scene text images. Our exploratory experiments, detailed in § 3, also underline the influential role of TP-HR in guiding the SR network to generate images with improved semantic accuracy for the model. Therefore, we introduce the TPEM to obtain the ETP, which can effectively guide the SR network, similarly to the efficacy of TP-HR. The core component of TPEM is the denoiser, denoted as $f_{\theta}$, which leverages the reverse diffusion process (Ho et al., 2020) to estimate the enhanced prior, providing substantial semantic guidance to the SR network.
3.2.1. Forward Diffusion Process
In the training phase, with a given HR image, denoted as $I^{\mathrm{hr}}\in\mathbb{R}^{2H\times 2W\times C}$, the TPG generates a sequence of recognition probabilities, referred to as $P^{\mathrm{h}}\in\mathbb{R}^{L\times|\mathcal{A}|}$, serving as our ground truth. $L$ is the length of the sequence and $|\mathcal{A}|$ is the cardinality of the recognizable letter set. Consequently, in line with Ho et al. (Ho et al., 2020), we incrementally introduce Gaussian noise denoted as $\epsilon$ to the initial variable $x_{0}=P^{\mathrm{h}}$ based on the timestamp, as follows:
(1) $q\left(x_{t}\mid x_{t-1}\right)=\mathcal{N}\left(x_{t};\sqrt{\alpha_{t}}x_{t-1},\left(1-\alpha_{t}\right)\boldsymbol{I}\right),$
Here, $\alpha_{t}$ is a hyperparameter that controls the variance of the added Gaussian noise at each time step. Leveraging the reparameterization trick (Kingma and Welling, 2014), we can express $x_{t}$ as:
(2) $x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,\quad\epsilon\sim\mathcal{N}(\mathbf{0},\boldsymbol{I}),$
where $\alpha_{i}\in[0,1]$, $\bar{\alpha}_{t}=\prod_{i=0}^{t}\alpha_{i}$, $t=1,2,\cdots,T$. As $T\to\infty$, $x_{T}$ converges to an isotropic Gaussian distribution. Consequently, during the inference phase, the forward diffusion process simplifies to initializing $x_{T}\sim\mathcal{N}(\mathbf{0},\boldsymbol{I})$.
3.2.2. Reverse Diffusion Process
In the reverse diffusion process, Gaussian noise gradually transforms into the ETP, denoted as $P^{\mathrm{e}}\in\mathbb{R}^{L\times|\mathcal{A}|}$, conditioned on the primary text prior $P^{\mathrm{l}}\in\mathbb{R}^{L\times|\mathcal{A}|}$. The latter is the output recognition probability sequence of the TPG, with $I^{\mathrm{lr}}$ as input. This process can be mathematically formulated as follows:
(3) $p_{\theta}\left(x_{t-1}\mid x_{t},P^{\text{l}}\right)=q\left(x_{t-1}\mid x_{t},f_{\theta}\left(x_{t},P^{\text{l}},t\right)\right),$
where $f_{\theta}$ is an MLP-based denoising network, the architecture of which can be found in the Supplementary Material. Similar to previous works (Liu et al., 2023; Xia et al., 2023; Yang et al., 2023; Zhao et al., 2023), we opt to directly estimate $P^{\mathrm{e}}$ instead of $\epsilon$ for performance improvement. Experiments in § 5 verifies the effectiveness of this design. The whole process is supervised by the MAE and CTC loss (Graves et al., 2006), given by:
(4) $\mathcal{L}_{\mathrm{diff}}=\lambda_{1}\underset{\mathcal{L}_{\mathrm{mae}}}{\underbrace{\left|\left|P^{\mathrm{h}}-P^{\mathrm{e}}\right|\right|_{1}}}+\lambda_{2}\mathcal{L}_{\mathrm{ctc}}^{\mathrm{t}},$
where $\lambda_{1}$ and $\lambda_{2}$ are the weight of the two losses. During the training phase, this reverse sampling process is a Markov process containing $T$ steps, which is computationally intensive. Therefore, in the inference phase, we adopt the sampling strategy of DDIM (Song et al., 2021) with $S$ steps ($S\ll T$) following previous works (Zhu et al., 2024a). Experiments in the Supplementary Material validate the effectiveness and efficiency of this design.
Besides, a feature alignment process between the shallow visual feature and the ETP is required to facilitate the SR process. We design a Feature Alignment Module (FAM) to obtain the aligned feature $F^{\mathrm{a}}\in\mathbb{R}^{H\times W\times C_{1}}$, where $C_{1}$ is the dimension of the shallow visual feature. The architecture of the FAM is shown in Figure 3(a).
3.3. Attention-Based Modulation Module
We design an Attention-based Modulation Module (AMM) to capture long-range dependence, thereby effectively restoring the visual structure of images with long or deformed text. As can be seen in Figure 2, AMM contains $N$ blocks, each block (i.e., $B_{i}$) including a simple convolutional layer, a Local Attention Module (LAM) and a Global Attention Module (GAM). For $B_{i}$, it receives the output of $B_{i-1}$, which is then concatenated with $F^{\mathrm{a}}$ channel-wisely to get $F_{i}^{\mathrm{c}}$, serving as the input of this block. Since the dimension of $F_{i}^{c}$ is located in $\mathbb{R}^{H\times W\times 2C_{1}}$, a convolutional layer is adopted to project it to the same space as $F_{i-1}^{\mathrm{o}}$, as:
(5) $F_{i}^{\mathrm{loc}}=\mathrm{Conv}\left(\mathrm{Concat}\left(F_{i-1}^{\mathrm{o}},F^{\mathrm{a}}\right)\right)\in\mathbb{R}^{H\times W\times C_{1}}.$
Firstly, a LAM is adopted to model the local similarity between intra- and inter-character features. Considering that for a scene text image, the horizontal contexts contain correlation information between characters while the vertical contexts contain internal features inside a character, such as the stroke information (Wang et al., 2020; Zhao et al., 2021), we propose to perform the strip-wise attention mechanism (Dong et al., 2022; Huang et al., 2023; Tsai et al., 2022) on the fusion feature to capture long-range dependence. Taking the horizontal attention as an example. As shown in Figure 3(b), $F_{i}^{\mathrm{loc}}$ is split into $W$ strips in the width dimension and the attention mechanism is applied to each of the $W$ strips. The resulting feature is then concatenated to form the horizontal feature $F_{i}^{\mathrm{loc(h)}}\in\mathbb{R}^{H\times W\times C_{1}}$. Likewise, the vertical attention mechanism can also result in the vertical feature $F_{i}^{\mathrm{loc(v)}}\in\mathbb{R}^{H\times W\times C_{1}}$. $F_{i}^{\mathrm{loc(h)}}$ and $F_{i}^{\mathrm{loc(v)}}$ are concatenated in the channel dimension and a convolutional layer is used to fuse them. Then the FFN with residual connection is introduced to perform non-linear transformation, given by:
(6) $F_{i}^{\mathrm{glo}}=F_{i}^{\mathrm{loc}}+\mathrm{Conv}\left(\mathrm{Concat}\left(F_{i}^{\mathrm{loc(h)}},F_{i}^{\mathrm{loc(v)}}\right)\right),$
(7) $\tilde{F}_{i}^{\mathrm{glo}}=F_{i}^{\mathrm{glo}}+\mathrm{FFN}\left(F_{i}^{\mathrm{glo}}\right)\in\mathbb{R}^{H\times W\times C_{1}}.$
However, the interaction between character-level features is limited in a local manner (Zhou et al., 2022). To encourage global interaction, we employ the strip-wise GAM as the complementation of the LAM, thus further modeling the global similarity between intra- and inter-character features. Specifically, the width dimension of $\tilde{F}_{i}^{\mathrm{glo}}$ is mer-ged with the channel dimension, leading to a feature map located in $\mathbb{R}^{H\times C_{1}W}$, as can be seen in Figure 3(c). The strip-wise attention mechanism is imposed on it to capture the global inter-character dependence, resulting in the feature $\tilde{F}_{i}^{\mathrm{glo(h)}}\in\mathbb{R}^{H\times C_{1}W}$. Likewise, the height dimension of $\tilde{F}_{i}^{\mathrm{glo}}$ is merged with the channel dimension, and the attention mechanism is applied to this feature map, leading to the feature $\tilde{F}_{i}^{\mathrm{glo(v)}}\in\mathbb{R}^{W\times C_{1}H}$. With the aid of them, the horizontal and vertical attention mechanisms can work globally to promote global-level interaction of character features. Similar workflow as Eq. (6) and (7) is adopted to get the final feature of $B_{i}$, namely $F_{i}^{\mathrm{o}}\in\mathbb{R}^{H\times W\times C_{1}}$.
Table 1. The recognition accuracy of some mainstream STISR methods on the three subsets of TextZoom. Best scores are bold.
  Methods Accuracy of ASTER (Shi et al., 2018) (%) Accuracy of MORAN (Luo et al., 2019) (%) Accuracy of CRNN (Shi et al., 2016a) (%)
Easy Medium Hard Average Easy Medium Hard Average Easy Medium Hard Average
LR 62.4 42.7 31.6 46.6 59.4 36.0 28.2 42.3 37.5 21.2 21.4 27.3
SRCNN (Dong et al., 2016) 69.4 43.4 32.2 49.5 63.2 39.0 30.2 45.3 38.7 21.6 20.9 27.7
SRResNet (Ledig et al., 2017) 69.6 47.6 34.3 51.3 60.7 42.9 32.6 46.3 39.7 27.6 22.7 30.6
RDN (Zhang et al., 2018) 70.0 47.0 34.0 51.5 61.7 42.0 31.6 46.1 41.6 24.4 23.5 30.5
LapSRN (Lai et al., 2017) 71.5 48.6 35.2 53.0 64.6 44.9 32.2 48.3 46.1 27.9 23.6 33.3
ESRT (Dong et al., 2016) 69.8 49.1 35.2 52.5 61.9 41.7 32.2 46.3 48.2 27.9 25.8 34.8
Omni-SR (Wang et al., 2023) 71.2 52.3 38.1 54.9 66.7 47.9 36.5 51.4 54.8 37.4 29.4 41.4
SRFormer (Zhou et al., 2023) 69.0 45.1 32.8 50.2 61.3 39.6 29.9 44.7 41.0 22.8 22.9 29.6
TSRN (Wang et al., 2020) 75.1 56.3 40.1 58.3 70.1 53.3 37.9 54.8 52.5 38.2 31.4 41.4
TBSRN (Chen et al., 2021) 75.7 59.9 41.6 60.1 74.1 57.0 40.8 58.4 59.6 47.1 35.3 48.1
PCAN (Zhao et al., 2021) 77.5 60.7 43.1 61.5 73.7 57.6 41.0 58.5 59.6 45.4 34.8 47.4
TG (Chen et al., 2022) 77.9 60.2 42.4 61.3 75.8 57.8 41.4 59.4 61.2 47.6 35.5 48.9
SGENet (TomyEnrique et al., 2024) 75.8 60.7 45.0 61.4 71.5 56.2 41.4 57.3 59.4 47.9 37.7 49.0
TPGSR (Ma et al., 2023) 78.9 62.7 44.5 62.8 74.9 60.5 44.1 60.5 63.1 52.0 38.6 51.8
TATT (Ma et al., 2022) 78.9 63.4 45.4 63.6 72.5 60.2 43.1 59.5 62.6 53.4 39.8 52.6
C3-STISR (Zhao et al., 2022) 79.1 63.3 46.8 64.1 74.2 61.0 43.2 60.5 65.2 53.6 39.8 53.7
TATT + DPMN (Zhu et al., 2023b) 79.3 64.1 45.2 63.9 73.3 61.5 43.9 60.4 64.4 54.2 39.2 53.4
TSAN (Zhu et al., 2023a) 79.6 64.1 45.3 64.1 78.4 61.3 45.1 62.7 64.6 53.3 38.8 53.0
TEAN (Shu et al., 2023) 80.4 64.5 45.6 64.6 76.8 60.8 43.4 61.4 63.7 52.5 38.1 52.2
MSPIE (Zhu et al., 2024b) 80.4 63.4 46.3 64.4 74.0 61.4 44.4 60.8 64.5 54.2 39.6 53.5
TCDM (Noguchi et al., 2024) 81.3 65.1 50.1 65.5 77.6 62.9 45.9 62.2 67.3 57.3 42.7 55.7
LEMMA (Guo et al., 2023) 81.1 66.3 47.4 66.0 77.7 64.4 44.6 63.2 67.1 58.8 40.6 56.3
RTSRN (Zhang et al., 2023) 80.4 66.1 49.1 66.2 77.1 63.3 46.5 63.2 67.0 59.2 42.6 57.0
RGDiffSR (Zhou et al., 2024) 81.1 65.4 49.1 66.2 78.6 62.1 45.4 63.1 67.6 56.5 42.7 56.4
TextDiff (Liu et al., 2023) 80.8 66.5 48.7 66.4 77.7 62.5 44.6 62.7 64.8 55.4 39.9 54.2
PEAN 84.5 71.4 52.9 70.6 79.4 67.0 49.1 66.1 68.9 60.2 45.9 59.0
HR 94.2 87.7 76.2 86.6 91.2 85.3 74.2 84.1 76.4 75.1 64.6 72.4
3.4. Multi-Task Learning
While previous researches (Guo et al., 2023; Ma et al., 2023, 2022; Zhang et al., 2023; Zhao et al., 2022) adopt the MTL paradigm by incorporating an additional text prior loss (Ma et al., 2022) to fine-tune the TPG for better text priors, our approach differs in the utilization of the MTL paradigm. We employ this paradigm not only to ensure image quality but also to facilitate text recognition. This distinction arises from the fact that STISR aims to simultaneously enhance the resolution and legibility of scene text images (Zhu et al., 2023b).
3.4.1. Image Restoration Task.
To begin with, we employ an image restoration task to generate high-quality SR images. As can be seen in Figure 3, the output of the last block of AMM, i.e., $F_{N}^{\mathrm{o}}$, is sent into a Super-Resolution Module (SRM) to get the SR image $I^{\mathrm{sr}}\in\mathbb{R}^{2H\times 2W\times C}$. It uses several convolutional layers with batch normalization and Mish (Misra, 2020) activation function for feature refinement and adopts a PixelShuffle (Shi et al., 2016b) operation to increase the resolution. To ensure the pixel-level and structure-level coherence between $I^{\mathrm{sr}}$ and $I^{\mathrm{hr}}$, the Mean-Square-Error (MSE) loss and the Stroke-Focused Module (SFM) loss (Chen et al., 2022) are applied between the two images, formulated as:
(8) $\mathcal{L}_{\mathrm{img}}=\lambda_{3}\underset{\mathcal{L}_{\mathrm{mse}}}{\underbrace{\left|\left|I^{\mathrm{hr}}-I^{\mathrm{sr}}\right|\right|_{2}^{2}}}+\lambda_{4}\underset{\mathcal{L}_{\mathrm{sfm}}}{\underbrace{\left|\left|A^{\mathrm{hr}}-A^{\mathrm{sr}}\right|\right|_{1}}},$
where $A$ is the attention map fetched from a Transformer-based recognizer (Chen et al., 2022). $\lambda_{3}$ and $\lambda_{4}$ are the weights of the MSE and SFM loss respectively.
3.4.2. Text Recognition Task.
Considering that high image quality may not be equal to superior recognition results (Guo et al., 2023; Liu et al., 2023), we introduce an Auxiliary Recognition Module (ARM), steering the model to generate SR images with promising legibility. In the inference phase, this module is abandoned, causing no extra computation complexity.
We exploit the classic and effective CNN-BiLSTM architecture applied in the STR task to design the ARM for simplicity (Shi et al., 2016a). Its output is a recognition probability sequence on the basis of $F_{N}^{\mathrm{o}}$ and we denote it as $P^{\mathrm{a}}\in\mathbb{R}^{L\times|\mathcal{A}|}$, where $L$ is the length of the sequence and $|\mathcal{A}|$ is the cardinality of the recognizable letter set. The CTC (Graves et al., 2006) loss denoted as $\mathcal{L}_{\mathrm{ctc}}^{\mathrm{a}}$ is imposed between $P^{\mathrm{a}}$ and the ground truth for better optimization. In a word, the total loss of the text recognition task is:
(9) $\mathcal{L}_{\mathrm{txt}}=\lambda_{5}\mathcal{L}_{\mathrm{ctc}}^{\mathrm{a}},$
where $\lambda_{5}$ is the weight of the CTC (Graves et al., 2006) loss for the text recognition task. In the training phase, we optimize the parameters of PEAN. The total loss is:
(10) $\mathcal{L}=\mathcal{L}_{\mathrm{diff}}+\mathcal{L}_{\mathrm{img}}+\mathcal{L}_{\mathrm{txt}}.$
4. Experiments
In this section, we first introduce the datasets and evaluation metrics. Then the implementation details are thoroughly described. Subsequently, comprehensive experiments are conducted to demonstrate that our proposed PEAN is an effective alternative for STISR.
4.1. Datasets and Evaluation Metrics
We conduct experiments on the TextZoom (Wang et al., 2020) benchmark. However, due to some inherent drawbacks of it, these experiments are not enough to reflect that PEAN surely has the ability to restore the complete visual structure. Therefore, we select 651 images with the resolution no greater than $16\times 64$ as LR images from the IIIT5K (Mishra et al., 2012), SVTP (Phan et al., 2013) and IC15 (Karatzas et al., 2015) datasets for evaluation. Details of the datasets and drawbacks of TextZoom can be found in the Supplementary Material. For TextZoom, following previous works (Wang et al., 2020; Chen et al., 2021, 2022; Ma et al., 2023, 2022; Zhao et al., 2022; Guo et al., 2023; Zhang et al., 2023), we adopt ASTER (Shi et al., 2018), MORAN (Luo et al., 2019) and CRNN (Shi et al., 2016a) for evaluation. In the Supplementary Material, we also adopt three recent Transformer-based recognizers, namely, MGP-STR (Wang et al., 2022), ABINet (Fang et al., 2021a) and VisionLAN (Wang et al., 2021) for evaluation. For the constructed dataset, PSNR and SSIM (Wang et al., 2004) are selected as metrics to evaluate the image quality.
4.2. Implementation Details
Our model is implemented with the Pytorch 1.10 deep learning library (Paszke et al., 2019). All of the experiments are conducted on 1 NVIDIA TITAN RTX GPU. In the training phase, the model is trained for 200 epochs and optimized by the AdamW (Loshchilov and Hutter, 2019) optimizer. The learning rate and the size of the mini-batch are set as 0.001 and 32 respectively. The weight of each loss is set as $\lambda_{1}=1$, $\lambda_{2}=1$, $\lambda_{3}=0.8$, $\lambda_{4}=75$, $\lambda_{5}=1$. In terms of the network architecture, the number of blocks in AMM is 6. The sampling timestep in the TPEM is set as 1. Following CRNN (Shi et al., 2016a), the number of convolutional layers applied in the ARM is 6. We first drop the TPEM and pre-train the model with TP-HR, then the TPEM is introduced and weights of parameters obtained by pre-training are initialized to continue the fine-tuning process of the model. Of note, if this setting is abandoned, PEAN can still achieve SOTA performance on the TextZoom (Wang et al., 2020) dataset. Ablation studies about the weight of loss functions and the pre-training and fine-tuning setting can be found in the Supplementary Material. In the main paper, we study our method on top of PARSeq (Bautista and Atienza, 2022) as the TPG. In the Supplementary Material, we also study the cases when CRNN (Shi et al., 2016a) and ABINet (Fang et al., 2021a) are adopted as the TPG.
Figure 4. Statistics on the performance of different text prior-based models with publicly available weights on images containing text of different lengths.
4.3. Comparing with State-of-the-Art Methods
We evaluate our proposed PEAN on the TextZoom benchmark (Wang et al., 2020) and compare its performance with several typical STISR methods. The results are presented in Table 1. It is evident that our proposed PEAN achieves new SOTA performance with significant improvements. For instance, when considering the recognition accuracy of ASTER (Shi et al., 2018), our model shows an average improvement of +4.2 compared with the existing SOTA model, namely, TextDiff (Liu et al., 2023). Furthermore, we provide statistics on the performance of different text prior-based models (Guo et al., 2023; Ma et al., 2022; Zhao et al., 2022) on images containing text of different lengths. For a fair comparison, all selected models use publicly available weights. As illustrated in Figure 4, PEAN outperforms other models across nearly every text length. It is particularly superior to previous works with images containing lengthy text, establishing itself as the first model capable of processing images containing text with up to 15 letters.
Figure 5. Visualization of SR images and their recognition results by ASTER. Red characters indicate wrong recognition results.
In addition, we provide some visualizations of samples from TextZoom recovered by several representative STISR models, as shown in Figure 5. With the assistance of the ETP, PEAN is able to generate SR images with improved semantic accuracy compared with previous text prior-based methods (Guo et al., 2023; Ma et al., 2022; Zhao et al., 2022). Although early methods such as TSRN (Wang et al., 2020) can generate SR images with correct semantic information for words like “cooking”, it is obvious that the visual structure of the text in images is disastrous.
Additionally, we perform experiments on the dataset we built without retraining the models. Improved performance on PSNR and SSIM in Table 2 shows that compared with previous representative STISR methods, PEAN is better at recovering the visual structure of the text in images.
Table 2. The performance of representative STISR models on the dataset we built.
  Methods PSNR SSIM
TSRN (Wang et al., 2020) 21.48 0.7743
TBSRN (Chen et al., 2021) 23.01 0.7967
TG (Chen et al., 2022) 21.84 0.7116
TATT (Ma et al., 2022) 23.31 0.8012
C3-STISR (Zhao et al., 2022) 21.03 0.7602
LEMMA (Guo et al., 2023) 22.09 0.7549
PEAN 24.24 0.8021
4.4. Ablation Study
Here we conduct ablation studies to demonstrate the effectiveness of components in our model. All the experiments are conducted on TextZoom and we report the recognition accuracy of ASTER (Shi et al., 2018).
4.4.1. Text Prior and the Enhancement Module
We conduct experiments to verify the importance of incorporating the text prior and the enhancement module into our model. The results shown in Table 3 justify that the introduction of the text prior can improve the performance of the SR process. However, this primary text prior from LR images is not robust enough to guide the SR network to generate SR images with high semantic accuracy. Considering that TP-HR can serve as a powerful alternative for guidance, we conduct an exploratory experiment wherein we substitute TP-LR with TP-HR. Results shown in the last row of the table demonstrate that leveraging the powerful prior information from HR images can significantly boost the performance of text prior-based STISR methods. This observation motivates us to design a module aimed at enhancing the primary text prior from LR images, thereby providing further guidance to the SR network for generating images with high semantic accuracy. Results displayed in the table indicate that our proposed TPEM is able to produce the ETP, which significantly improves the performance over the model with TP-LR among all subsets by +6.8 on average.
Table 3. Analysis of the impact of the text prior and the TPEM. The last row shows the results of the exploratory experiment as illustrated in Figure 1(f).
  TP-LR TPEM TP-HR Easy Medium Hard Average
75.7 60.2 42.1 60.4
79.7 62.3 46.1 63.8
84.5 71.4 52.9 70.6
88.4 75.5 61.3 75.9
Since previous works such as Zhao et al. (Zhao et al., 2022) introduce a Language Model (LM) (Fang et al., 2021a) to rectify the text prior, here we compare the TPEM with several LMs used in other tasks of scene text images. Among them, GSRM is an autoregressive-based LM proposed by SRN (Yu et al., 2020), while PD is the parallel Transformer decoder in PIMNet (Qiao et al., 2021) based on an easy-first decoding strategy. BCN (Fang et al., 2021a) is the autoencoding-based LM applied in C3-STISR (Zhao et al., 2022), proposed in ABINet (Fang et al., 2021a). Experimental results are shown in Table 4, from which we can conclude that the employment of the diffusion-based TPEM brings more performance gain compared with other variants. This can be attributed to the powerful distribution mapping capability (Xia et al., 2023) of diffusion models.
Table 4. Ablation study on different enhancement modules.
  Methods Easy Medium Hard Average
GSRM (Yu et al., 2020) 76.3 58.8 40.7 59.7
PD (Qiao et al., 2021) 78.1 61.6 43.4 62.1
BCN (Fang et al., 2021a) 79.6 61.9 44.8 63.2
PEAN 84.5 71.4 52.9 70.6
4.4.2. Optimization Objective for Training the TPEM
Regarding to most of the diffusion-based models (Saharia et al., 2023; Gao et al., 2023a; Ren et al., 2023; Rombach et al., 2022), researchers typically constrain the estimated noise (denoted as $\epsilon$) through the loss function, as supported by Ho et al. (Ho et al., 2020), for the sake of improving sample quality. However, as discussed in § 3.2.2, in the context of the TPEM, we choose to directly estimate the ETP (denoted as $P^{\mathrm{e}}$) instead of focusing on the noise for enhanced performance. In this regard, we conduct experiments with both optimization objectives, and the results are presented in Table 5. We argue that selecting $P^{\mathrm{e}}$ as the optimization target can yield superior performance in our proposed PEAN because it encourages the network to concentrate on the enhancement of the text prior itself. This approach also opens the possibility of introducing additional constraints on the ETP, such as the CTC loss (Graves et al., 2006) denoted as $\mathcal{L}_{\mathrm{ctc}}^{\mathrm{t}}$.
Table 5. Ablation study about the objective of optimization.
  Objectives Easy Medium Hard Average
$\epsilon$ 80.0 65.6 46.3 65.0
$P^{\mathrm{e}}$ 84.5 71.4 52.9 70.6
4.4.3. Impact of the AMM
Here we perform experiments to justify the superiority of the AMM adopted in our model. As presented in Table 6, the comparison reveals the following points: (1) Our proposed AMM handles the STISR task in a more effective way than the CNN-BiLSTM-based SRB (Wang et al., 2020). The local inductive bias of the CNN (Dosovitskiy et al., 2020; Raghu et al., 2021) and the rigid nature of BiLSTM (Vaswani et al., 2017; Zhu et al., 2023b) limit its performance in recovering the visual structure of images with long or deformed text (Chen et al., 2021; Ma et al., 2022), while the AMM can solve this problem. (2) The typical ViT-based architectures (Dong et al., 2022; Dosovitskiy et al., 2020; Liu et al., 2021) show trivial performance, as a result of the tendentious design for high-level vision tasks, which neglects the inherent characteristic of STISR. (3) Although employing the strip-wise attention mechanism, Stripformer (Tsai et al., 2022), which also leverages the conditional positional encodings (Chu et al., 2023) and several residual blocks for the image deblurring task, fails in the STISR task. On average, PEAN obtains an improvement of the performance of +5.9 from the AMM. Further experiments in the Supplementary Material show that even without the MTL paradigm, the AMM still outperforms the SRB (Wang et al., 2020).
Table 6. Ablation study on the impact of the AMM.
  Methods Easy Medium Hard Average
SRB (Wang et al., 2020) 80.1 64.4 46.4 64.7
ViT (Dosovitskiy et al., 2020) 81.8 65.7 49.5 66.7
Swin (Liu et al., 2021) 73.8 55.1 39.0 57.1
CSWin (Dong et al., 2022) 70.2 52.9 37.2 54.5
Stripformer (Tsai et al., 2022) 72.9 53.6 37.3 55.7
AMM 84.5 71.4 52.9 70.6
4.4.4. Effect of the MTL Paradigm
In this part, we conduct experiments to show the effect of the MTL paradigm introduced in our work. The results shown in Table 7 indicate that: (1) The stroke-based SFM loss (Chen et al., 2022) provides an average contribution of +4.6 for the improved performance, which convinces us that the preservation of visual structure plays a vital role in STISR. (2) The text recognition task employed in our work is also indispensable. The CTC (Graves et al., 2006) loss applied on the output of ARM ($\,\mathcal{L}_{\mathrm{ctc}}^{\mathrm{a}}$) brings correct semantic constraint for the AMM. Besides, the CTC loss imposed on the ETP ($\,\mathcal{L}_{\mathrm{ctc}}^{\mathrm{t}}$) ensures the coherence between the text prior and the ground truth text label, guiding the training of the enhancement module in a precise way. We provide more detailed experiments in the Supplementary Material.
Table 7. Ablation study on the effect of MTL.
  Loss Functions Easy Medium Hard Average
$\mathcal{L}_{\mathrm{mse}}$ 76.2 58.8 41.5 59.9
+$\,\mathcal{L}_{\mathrm{sfm}}$ 79.2 64.3 47.0 64.5
+$\,\mathcal{L}_{\mathrm{mae}}$ 79.6 65.1 47.1 64.9
+$\,\mathcal{L}_{\mathrm{ctc}}^{\mathrm{t}}$ 81.4 68.8 50.7 67.9
+$\,\mathcal{L}_{\mathrm{ctc}}^{\mathrm{a}}$ 84.5 71.4 52.9 70.6
4.5. Representation Analysis
In this section, we delve into the reasons about the ability of the ETP that guides the SR network in generating images with improved semantic accuracy. In Table 3, we observe superior performance for PEAN with TP-HR, while PEAN with TP-LR exhibits poorer results compared to PEAN with ETP, as demonstrated in Table 3. To further investigate these observations, following the common analysis setting on Vision Transformers (Park and Kim, 2022; Raghu et al., 2021), we employ the linear centered kernel alignment (CKA) (Kornblith et al., 2019) to assess the similarity of representations among these three models.
[…]
Consequently, we can conclude that the power of the ETP lies in its ability to make the representations learned by AMM and SRM more similar to those learned by the corresponding modules in PEAN (w/ TP-HR), which is known for its superior performance. This study further justifies that the combination of our proposed TPEM and AMM brings out powerful capabilities.
Figure 6. Results of representation comparison by CKA (Kornblith et al., 2019).
5. Conclusion
In this paper, we propose a Prior-Enhanced Attention Network (PEAN) for scene text image super-resolution (STISR). Specifically, we design a Text Prior Enhancement Module (TPEM) to provide the ETP for the subsequent SR process, enabling SR images to contain accurate semantic information. Moreover, an Attention-based Modulation Module (AMM) is devised to obtain local and global coherence in scene text images, which can recover the visual structure of images with text in various sizes and deformations. Additionally, we introduce the Multi-Task Learning (MTL) paradigm to improve the legibility of LR images. Experiments demonstrate that our proposed PEAN achieves SOTA performance through the interaction of these designs. We believe our work will serve as a strong baseline for future works, and will push forward the research of STISR as well as other sub-fields of scene text images.